A machine will reduce labor costs by $3,000 per month. It has a lifespan of 2 years, and no salvage value. Maintenance costs start at $50 per month, and increase by $50 each month. (So for months 1,2,3,4 costs are $50, $100, $150, $200.) Assuming a nominal annual discount rate of 6%, what’s the most you should pay for such a machine? (Neglect tax and depreciation.) (You should use the tables for this for EIT practice.)
Answer #1

i = 6% / 12 = 0.5% per month

t = 2 * 12 = 24 months

Present value of savings and maint cost = 3000 * (P/A,0.5%,24) -50 * (P/G,0.5%,25)*(F/P,0.5%,1)

= 3000 * 22.562866 - 50 * 275.268557*1.005

= 53856.35

Maximum value that can be paid = 53856 (Nearest Dollar)